A new car is purchased for 22300 dollars. The value of the car depreciates at 10.25% per year. What will the value of the car be, to the nearest cent, after 6 years?



1
Expert's answer
2021-05-13T08:18:13-0400

Principal=$22300

Time=6 years

Rate of depreciation=10.25%

Formula used "A=P(1-\frac{R}{100})^{t}"

Where

A=amount

P=principal

t=time

R=rate of depreciation

"A=22300(1−\frac{10.25}{100})^{6}=11654.98"

Therefore the value of the car will be $11654.98